OpenStudy (anonymous):

Determine the coordinates of a unit vector that is perpendicular to both a = <1, 1, 1> and b = <1, 3 -1> and makes an obtuse angle with the z-axis.

OpenStudy (anonymous):

I calculated the cross-product- it's <-4, 2, 2> and calculated the unit vector of that, but the answer is wrong. I think it's because of the obtuse thing- is there a specific formula I should be using?

OpenStudy (anonymous):

We can represent the \(z\) axis with the vector \(\langle 0,0,1\rangle\).

OpenStudy (anonymous):

What I mean is, such is a vector that is parallel to the positive \(z\) axis.

OpenStudy (anonymous):

So I find a vector parallel to the z-axis, but also perpendicular to a and b?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The reason I did this is because you can find the angle between to vectors.

OpenStudy (anonymous):

two vectors

OpenStudy (anonymous):

\[ |a||b|\cos\theta = |a\cdot b| \]

OpenStudy (anonymous):

I think the question is a bit silly though, because it will always form an acute angle with the \(z\) axis if you count the positive and negative... so I assume they're only considering the positive part of the axis.

OpenStudy (anonymous):

 In general if we project onto the \(xy\) plane we can see that our angle will be a right angle.

OpenStudy (anonymous):

Any vector with a positive \(z\) value will be acute to the positive \(z\) axis, and any vector with a negative \(z\) value will be obtuse to the positive \(z\) axis.

OpenStudy (anonymous):

Also... I should have put:\[ |a||b|\cos\theta = a\cdot b \]

OpenStudy (anonymous):

Had we let \(b=\langle 0,0,1\rangle\), then we would have \(|b| = 1\) and \(a\cdot b = a_z\)

OpenStudy (anonymous):

I didn't realize that, ha ha. Um, so I'm sorry, but I'm still stuck- are there mutiple vectors that are perpendicular to two vectors? Because the only one I can find is <-4, 2, 2>, and that does not result in an obtuse angle with the positive z-axis.

OpenStudy (anonymous):

Consider that if \(a\cdot b = 0\), then \((-a)\cdot b = -a\cdot b = 0\).

OpenStudy (anonymous):

We can always say that in \(\mathbb R^3\) there will be two vectors orthogonal to any pair of vectors.

OpenStudy (anonymous):

\[ u\times v = -(v\times u) \]

OpenStudy (anonymous):

So, if you did the cross of \(a\) and \(b\), maybe try \(b\) and \(a\). It's not a commutative operation.

OpenStudy (anonymous):

Alright, thanks, this one is obtuse. Unit vector calculation is: \[unit vector =\frac{ v }{ |v| }\] right? just making sure.

OpenStudy (anonymous):

Alright, got the answer! Thanks for your help!